Normally it takes John 2.5 hours to drive from his house to his office. Due to yesterday's snowy weather he was driving 10 mph slower than usually, and it took 3 hours. How many miles is the office away from John's house?
Answer by ikleyn(52765)

.
Let x be the regular average rate, in miles per hour.


Under the severe condition, the average rate was  (d-10) miles per hour.


The distance is the same:


2.5*x = 3*(x-10).


Simplify and solve:


2.5x = 3x - 30

30 = 3x - 2.5x

30 = 0.5x  ===================>  x = 30/0.5 = 60 milies per hour.


ANSWER.  The distance is  60*2.5 = 150 miles.

Solved.


/\/\/\/\/\/\/\/


I want to remind you that as an educated visitor, you MUST express your thanks to EACH the solution
that I produce in response to your request.


I forced to make this reminder because MANY (if not all) visitors IGNORE THIS RULE. Unfortunately.

Which I consider as an extreme degree of discourtesy.